a sailor has a rope that is 72feet long.It is cut into two pieces with one piece three-fifths as long as the other.find the lenght of each piece of rope.
Answer by SuperGirl98734(12)

if...
piece 1 = a
piece 2 = (a+3/5)
then...
a+(a+3/5)=72
-3/5
a+a=71 2/5
2a=71 2/5
/2
a=35 7/10
so...
piece 1 = 35 7/ 10
piece 2 = 36 3/10
hope that was helpful and easy to understand :)